# Natural Language Processing in Patents: A Survey

### v2

@@ Abstract
Patents, encapsulating crucial technical and legal information, present a rich domain for natural language processing (NLP) applications. As NLP technologies evolve, large language models (LLMs) have demonstrated outstanding capabilities in general text processing and generation tasks. However, the application of LLMs in the patent domain remains under-explored and under-developed due to the complexity of patent processing. Understanding the unique characteristics of patent documents and related research in the patent domain becomes essential for researchers to apply these tools effectively. Therefore, this paper aims to equip NLP researchers with the essential knowledge to navigate this complex domain efficiently. We introduce the relevant fundamental aspects of patents to provide solid background information, particularly for readers unfamiliar with the patent system. In addition, we systematically break down the structural and linguistic characteristics unique to patents and map out how NLP can be leveraged for patent analysis and generation. Moreover, we demonstrate the spectrum of text-based patent-related tasks, including nine patent analysis and four patent generation tasks.
Natural Language Processing in Patents: A Survey
Lekang Jiang  and Stephan Goetz University of Cambridge {lj408, smg84}@cam.ac.uk
@@ 1 Introduction
Researchers have investigated machine learning (ML) and natural language processing (NLP) methods for the patent field with highly technical and legal texts (Krestel et al., 2021). In addition, the recent large language models (LLMs) have demonstrated outstanding capabilities across a wide range of general domain tasks (Zhao et al., 2023; Min et al., 2023). These models are promising to become valuable tools in managing and drafting patent literature, the crucial resource that documents technological advances.
However, compared to the significant success of LLMs in the general domain, the application of LLMs in patent-related tasks remains under-explored due to the texts’ and the field’s complexity. NLP researchers need to deeply understand the unique characteristics of patent documents to develop useful models for the patent field. Therefore, we aim to equip NLP researchers with the essential knowledge by presenting this highly auspicious but widely neglected field to the NLP community.
@@
Previous surveys reported the early stages of smart and automated methods for patent analysis (Abbas et al., 2014), first deep learning methods, which opened a wider range of still simpler patent tasks (Krestel et al., 2021), or specific individual patent tasks, such as patent retrieval (Shalaby and Zadrozny, 2019). The recent advancements in NLP were unforeseen, particularly the performance boost when language models are scaled up. Accordingly, we specifically delineate a survey of popular NLP methodologies for patent documents with a special focus on the most recent and evolving techniques. We have included the two applications of patent analysis and generative patent tasks (Fig. 1). Whereas analysis focuses on understanding and usage of individual patent documents or a group of patents, generation tasks aim at automatically generating patent texts.
@@
    We analyze the unique structural and linguistic characteristics of patent texts with NLP techniques to process these texts in Section 3. Readers will readily understand the challenges of automated patent processing and the development trends of NLP in the patent field.
@@
    We present a detailed synthesis of patent data sources, alongside tailored datasets specifically designed for different patent tasks in Section 4. By providing these ready-to-use datasets, we aim to eliminate the extensive time and effort for data preparation in the patent domain.
@@
    We evaluate nine patent analysis tasks (subject classification, patent retrieval, information extraction, novelty prediction, granting prediction, litigation prediction, valuation, technology forecasting, and innovation recommendation) in Section 5 and four generation tasks (summarization, translation, simplification, and patent writing) in Section 6. Specifically, we systematically demonstrate task definitions and relevant methodologies in detail. We show a comprehensive yet accessible insight into this area, enabling readers to grasp the nuances of the field more effectively.
@@
    We identify current challenges and point out potential future research directions in Section 7. By highlighting these areas, we hope to encourage further research and development in automated patent tasks, leading to more efficient and effective methods in the future.
@@ 2.1 Patent Document
Patent documents are central elements for the protection of intellectual property and further document inventions. Patents require applicants and/or inventors to publicly disclose their inventions in detail to secure exclusive rights and obtain benefits in return. Patent documents describe new inventions and delineate the scope of patent rights granted to patent holders. These documents are key parts of the patenting process and are publicly accessible typically 18 months after the application or first filing date. The format and content can vary by jurisdiction but normally include the following elements. Fig. 2 displays an example patent document.
@@ 2.2 Patent Life Cycle
Pre-grant Stage. In the beginning, inventors conceptualize, design, and develop their inventions. If inventors hope to obtain patent protection for their inventions, they need to apply to the patent office. To ensure novelty and inventiveness (non-obviousness), inventors may preventively search for existing patents and public disclosures to avoid unnecessary cost and effort in case of prior disclosures. Additionally, patent documents need drafting, describing the invention in detail, including its specifications, claims, abstract, and accompanying drawings. The drafting process typically requires the expertise of a patent professional, such as a patent engineer, attorney, or agent. After the patent application is submitted to the patent office, examiners at the office will screen and evaluate the application for compliance with formal, legal, and technical requirements. This examination typically involves correspondence of the examiner with the inventor or their representation, where clarifications, amendments, or arguments are submitted. [^1] Amendments must not introduce novel features, i.e., inventive content, to avoid losing the original date stamp or the entire file. Notably, an assessment will determine if the invention meets the criteria for patentability, including novelty, inventiveness (non-obviousness), and typically less strictly commercial utility. The requirement of utility may be less strict as offices may see that as the applicant’s problem, except for certain cases. The substantive examination compares the invention with similar documents from the patent literature or any other public source dated earlier, which were found in an initial search by the office. Upon examination and potential resolution of any objections, the patent is either granted—rarely in its original, more frequently in a restricted form based on the identified prior art—or rejected. If the original priority date should be maintained, the invention must not be extended during the examination. [^2] Such additional material should not be added casually but would require proper announcement with a new filing that claims the priority of the base filing. Otherwise, the inadmissibly extended file may be rejected or later challenged in the granted stage.
@@ 2.3 Related Surveys
Patent Analysis. Various surveys in the past summarized certain aspects of patent analysis from a knowledge or procedural perspective. The work of Abbas et al. (2014) represents early research on patent analysis. These early methods included text mining and visualization approaches, which paved the way for future research. Deep learning for knowledge and patent management started more recently. Krestel et al. (2021) identified eight types of sub-tasks attractive for deep learning methods, specifically supporting tasks, patent classification, patent retrieval, patent valuation, technology forecasting, patent text generation, litigation analysis, and computer vision tasks. Some surveys reviewed specific topics of patent analysis. For example, Shalaby and Zadrozny (2019) surveyed patent retrieval, i.e., the search for relevant patent documents, which may appear similar to a web search but has substantially different objectives as well as constraints and has a legally-defined formal character.
Patent Usage. The patent literature records major technological progress and constitutes a knowledge database. It is well known that the translation of methods from one domain to another or the combination of technology from various fields can lead to major innovation. Thus, the contents of the patent literature appear highly attractive for systematic analysis and innovation design. However, the language of modern patents has substantially evolved and diverged from normal technical writing. Recent patent documents are typically hardly digestible for the normal reader and also contain deceptive elements added for increasing the scope or camouflage important aspects. Appropriate data science techniques need to consider such aspects to mine patent databases for engineering design (Jiang et al., 2022b). Data science applied to this body of codified technical knowledge can not only inform design theory and methodology but also form design tools and strategies. Particularly in early stages of innovation, language processing can help with ideation, trend forecasting, or problem–solution matching (Just, 2024).
@@ 3.1 Long Context
Most research so far has focused on short texts, such as patent abstracts and claims. However, titles and abstracts of patents can be surprisingly generic. Therefore, using patent descriptions that provide comprehensive details and specific embodiment of the invention for patent analysis is highly important, which is neglected by most of the current research. A possible reason is that previous language models cannot handle such long inputs. According to a recently proposed patent dataset (Suzgun et al., 2023), the average number of tokens of a patent description exceeds 11,000, which is longer than the context limit of many previous language models. For example, Llama-2 supports the context length of 4,000 tokens (Touvron et al., 2023). The long context of patent descriptions causes critical challenges for patent analysis.
Notably, researchers have investigated increasing or otherwise handling the context length of LLMs (Chen et al., 2023; Jiang et al., 2023a; Xiong et al., 2023). Xiong et al. (2023), for example, introduced a series of long-context LLMs, which can support the context windows up to 32,768 tokens. Moreover, the state-of-the-art GPT-4 model [^3] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo and Llama-3.1444https://ai.meta.com/blog/meta-llama-3-1/?_fb_noscript=1 even support a context length of 128,000 tokens, which appears promising to process the long patent descriptions. Other LLMs also showed a trend towards long context capabilities, such as Falcon-180B (Almazrouei et al., 2023), Gemini 1.5555https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ and Claude 3.5666https://www.anthropic.com/news/claude-3-5-sonnet.
@@ 3.4 LLMs for Patent Processing
Patent texts are distinct from everyday texts with respect to long context length, in-depth technical complexity, and high precision to ensure the patent can be granted, defended, and enforced. This type of language often necessitates specific training and experience in patents for accurate reading and interpretation. However, humans usually struggle with names and phrases that conflict with their prior knowledge, or numerous redefined or otherwise complex terms or structures. In contrast, LLMs based on specified training should in principle handle complex structures and newly coined terms that deviate from everyday language, including the specialized vocabulary of experts in the field. Therefore, the recent advancements in LLMs for generative tasks and language processing appear ideally suited for the patent literature.
@@
Despite the apparent fundamental compatibility of LLMs for knowledge extraction and language processing, the application of LLMs in the patent domain remains underdeveloped and not yet highly prominent. Previous studies used word embeddings (e.g., Word2Vec (Mikolov et al., 2013)) and deep learning models (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) for patent analysis tasks. As transformers (Vaswani et al., 2017) showed significant potential in text processing, researchers started to develop transformer-based language models, such as BERT (Devlin et al., 2018) and GPT (Radford et al., 2018). So far, most LLMs adapted and tested in the patent domain are on the level of GPT-2 (Radford et al., 2019) with approximately 1.5 billion parameters. The recent large-sized models with outstanding capabilities have not been extensively investigated in the patent field.
@@
Previous research efforts have included patent analysis, data extraction, and automation of procedures. However, the lack of benchmark tests, such as reference datasets and established metrics, hinders performance evaluation and comparison across different methods. The effectiveness of LLMs depends on the quality of training data. To this end, we have compiled sources and databases of patents with curated datasets tailored for various patent-related tasks in Section 4. Although patent offices have released raw documents for years, publicly available datasets for specific tasks remain scarce. Numerous studies continue to use closed-source data for training and evaluation. Furthermore, patent offices do not provide pre-processed data or broad access to well-structured documents from the process around patents. Although in many countries, patent documents as the disclosure of inventions are considered public domain, they offer only the manual review of individual documents. [^8] This aspect may reflect an important discussion to be led by the field and society. As in other domains, there may be a missing or unclear consensus on which data can be used for machine learning. In stark contrast to other domains, where even books and copyrighted materials may be processed in large numbers, sometimes to create new works that compete with the originals, the patent system is a bit different. The entire modern patent system was developed and is justified as a social contract, trading invention disclosure for temporary exclusion from competition. The descriptions are often public domain and related correspondence with patent offices is made public to ensure transparency. Thus, imposing paywalls would typically violate the mission of patent offices. Admittedly, big-data methods and artificial intelligence were not considerations during the establishment of these legal frameworks. Patent offices may still restrict or complicate larger-scale data access to parts or all contents to avoid commercial exploitation. Some patent offices offer costly paid database access as a stream of revenue, which can be prohibitively expensive for academic research. Important context to know is that in many places patent offices need to fund their own operating expenses (United States Patent and Trademark Office, 2023; President of the European Patent Office and Administrative Council of the European Patent Organisation, 2021; van Pottelsberghe de la Potterie and François, 2009; Frakes and Wasserman, 2014). Therefore, society—which the patent offices are ultimately meant to serve—must reach a consensus on how to manage data access for commercial versus academic purposes, potentially leading to both proprietary and open-source models.

### v3

# Natural Language Processing in the Patent Domain: A Survey

@@ Abstract
Patents, which encapsulate crucial technical and legal information in text form and referenced drawings, present a rich domain for natural language processing (NLP) applications. As NLP technologies evolve, large language models (LLMs) have demonstrated outstanding capabilities in general text processing and generation tasks. However, the application of LLMs in the patent domain remains under-explored and under-developed due to the complexity of patents, particularly their language and legal framework. Understanding the unique characteristics of patent documents and related research in the patent domain becomes essential for researchers to apply these tools effectively. Therefore, this paper aims to equip NLP researchers with the essential knowledge to navigate this complex domain efficiently. We introduce the relevant fundamental aspects of patents to provide solid background information. In addition, we systematically break down the structural and linguistic characteristics unique to patents and map out how NLP can be leveraged for patent analysis and generation. Moreover, we demonstrate the spectrum of text-based and multimodal patent-related tasks, including nine patent analysis and four patent generation tasks. [^1] This is an updated version of the preprint. The final version is published in Artificial Intelligence Review. Please refer to and cite the final version: https://doi.org/10.1007/s10462-025-11168-z.
Natural Language Processing in the Patent Domain: A Survey
Lekang Jiang  and Stephan M. Goetz University of Cambridge {lj408, smg84}@cam.ac.uk
@@ 1 Introduction
Researchers have investigated machine learning (ML) and natural language processing (NLP) methods for the patent field with highly technical and legal texts (Krestel et al., 2021). In addition, the recent large language models (LLMs) have demonstrated outstanding capabilities across a wide range of general domain tasks (Zhao et al., 2023; Min et al., 2023). Moreover, the expansion of the latest general LLMs with a graphical component to form multimodal models (Huang et al., 2024) may further enhance the capabilities in processing patents, which include text and drawings. These models are promising to become valuable tools in managing and drafting patent literature, the crucial resource that documents technological advances.
However, compared to the significant success of LLMs in the general domain, the application of LLMs in patent-related tasks remains under-explored due to the texts’ and the field’s complexity. NLP and multimodal model researchers need to deeply understand the unique characteristics of patent documents to develop useful models for the patent field. Therefore, we aim to equip researchers with the essential knowledge by presenting this highly auspicious but widely neglected field to the NLP community.
@@
Previous surveys reported the early stages of smart and automated methods for patent analysis (Abbas et al., 2014), first deep learning methods, which opened a wider range of still simpler patent tasks (Krestel et al., 2021), or specific individual patent tasks, such as patent retrieval (Shalaby and Zadrozny, 2019). The recent advancements in language and multimodal models were unforeseen, particularly the performance boost when models are massively scaled up (Kaplan et al., 2020). Accordingly, we specifically delineate a survey of popular methodologies for patent documents with a special focus on the most recent and evolving techniques. We have included the two applications of patent analysis and generative patent tasks (Fig. 1). Whereas analysis focuses on understanding and usage of individual patent documents or a group of patents, generation tasks aim at automatically generating patent texts.
@@
    We analyze the unique structural and linguistic characteristics of patent texts with language processing and multimodal techniques to process these texts in Section 3. Readers will readily understand the challenges of automated patent processing and the development trends of NLP in the patent field.
@@
    We present a detailed synthesis of patent data sources, alongside tailored datasets specifically designed for different patent tasks, in Section 4. With the collection of these ready-to-use datasets, we aim to eliminate the extensive time and effort for data preparation in the patent domain.
@@
    We evaluate nine patent analysis tasks (subject classification, patent retrieval, information extraction, novelty prediction, granting prediction, litigation prediction, valuation, technology forecasting, and innovation recommendation) in Section 5 and four generation tasks (summarization, translation, simplification, and patent writing) in Section 6. Specifically, we systematically demonstrate task definitions and relevant methodologies in detail. We show a comprehensive yet accessible insight into this area, which should enable readers to grasp the nuances of the field more effectively.
@@
    We identify current challenges and point out potential future research directions in Section 7. By highlighting these areas, we hope to encourage further research and development in automated patent tasks to stimulate more efficient and effective methods in the future.
@@ 2.1 Patent Document
Patent documents are central elements for the protection of intellectual property and also document inventions. Patents require applicants and/or inventors to publicly disclose their inventions in detail to secure exclusive rights and obtain benefits in return. Patent documents describe new inventions and delineate the scope of patent rights granted to patent holders. These documents are key parts of the patenting process and are publicly accessible typically 18 months after the application or the first filing date. The format and content can vary by jurisdiction but normally include the following elements. Fig. 2 displays an example patent document.
@@ 2.2 Patent Life Cycle
Pre-Grant Stage. In the beginning, inventors conceptualize, design, and develop their invention. If inventors hope to obtain patent protection for their invention, they need to apply to the patent office. To ensure novelty and inventiveness (non-obviousness), inventors may preventively search for existing patents and public disclosures to avoid unnecessary cost and effort in case of prior disclosures. Additionally, patent documents need drafting to describe the invention in detail, including its specifications, claims, abstract, and accompanying drawings. The drafting process typically requires the expertise of a patent professional, such as a patent engineer, attorney, or agent. After the patent application is submitted to the patent office, examiners at the office will screen and evaluate the application for compliance with formal, legal, and technical requirements. This examination typically involves correspondence of the examiner with the inventor or their representation, where clarifications, amendments, or arguments are submitted. [^2] Amendments must not introduce novel features, i.e., inventive content, to avoid losing the original date stamp or the entire file. Notably, an assessment will determine if the invention meets the criteria for patentability, including novelty, inventiveness (non-obviousness), and typically less strictly commercial utility. The requirement of utility may be less strict as offices may see that as the applicant’s problem, except for certain cases. The substantive examination compares the invention with similar documents from the patent literature or any other public source dated earlier, which were found in an initial search by the office. Upon examination and potential resolution of any objections, the patent is either granted—rarely in its original, more frequently in a restricted form based on the identified prior art—or rejected. If the original priority date should be maintained, the invention must not be extended during the examination. [^3] Such additional material should not be added casually but would require proper announcement with a new filing that typically claims the priority of the base filing for all aspects already described in the latter. Otherwise, the inadmissibly extended file may be rejected or later challenged in the granted stage.
@@ 2.3 Related Surveys
Patent Analysis. Various surveys in the past summarized certain aspects of patent analysis from a knowledge or procedural perspective. The work of Abbas et al. (2014) represents early research on patent analysis. These early methods included text mining and visualization approaches, which paved the way for future research. Deep learning for knowledge and patent management started more recently. Krestel et al. (2021) identified eight types of sub-tasks attractive for deep learning methods, specifically supporting tasks, patent classification, patent retrieval, patent valuation, technology forecasting, patent text generation, litigation analysis, and computer vision tasks. Some surveys reviewed specific topics of patent analysis. For example, Shalaby and Zadrozny (2019) surveyed patent retrieval, i.e., the search for relevant patent documents, which may appear similar to a web search but has substantially different objectives as well as constraints and has a legally-defined formal character. A concurrent survey introduced the latest development of patent retrieval and posed some challenges for future work (Ali et al., 2024). In addition, Balsmeier et al. (2018) highlighted how machine learning and NLP tools could be applied to patent data for tasks such as innovation assessment. It focused more on methodological advancement in patent data analysis that complements the broader survey of tools and models used in NLP for patent processing.
Patent Usage. The patent literature records major technological progress and constitutes a knowledge database. It is well known that the translation of methods from one domain to another or the combination of technology from various fields can lead to major innovation. Thus, the contents of the patent literature appear highly attractive for systematic analysis and innovation design. However, the language of modern patents has substantially evolved and diverged from normal technical writing. Recent patent documents are typically hardly digestible for the normal reader and also contain deceptive elements added for increasing the scope or to camouflage important aspects. Appropriate data science techniques need to consider such aspects to mine patent databases for engineering design (Jiang et al., 2022b). Data science applied to this body of codified technical knowledge cannot only inform design theory and methodology but also form design tools and strategies. Particularly in early stages of innovation, language processing can help with ideation, trend forecasting, or problem–solution matching (Just, 2024).
@@ 3.1 Long Context
Most research so far has focused on short texts, such as patent abstracts and claims. However, titles and abstracts of patents can be surprisingly generic. Therefore, using patent descriptions that provide comprehensive details and specific embodiment of the invention for patent analysis is highly important, which is neglected by most of the current research. A possible reason is that previous language models cannot handle such long inputs. According to a recently proposed patent dataset (Suzgun et al., 2023), the average number of tokens of a patent description exceeds 11,000, which is longer than the context limit of many previous language models. For example, Llama-2 supports the context length of 4,000 tokens (Touvron et al., 2023). The reason for the limited context length of many language models is the rapid growth of computational complexity associated with self-attention. It grows because the number of attention relationships increases with the context length. As self-attention often considers every pair of tokens, the computational complexity can grow with the context length squared ($\mathcal{O}(n^{2})$). The long context of patent descriptions causes critical challenges for patent analysis.
Notably, researchers have investigated increasing or otherwise handling the context length of LLMs (Chen et al., 2023; Jiang et al., 2023a; Xiong et al., 2023). Xiong et al. (2023), for example, introduced a series of long-context LLMs, which can support the context windows up to 32,768 tokens. A mix of reduced short-range and long-range attention instead of exhaustive attention can reduce the computational and memory burden (Kovaleva et al., 2019; Beltagy et al., 2020). Moreover, the latest very large models, such as GPT-$4$https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo (OpenAI, 2023) and Llama-3.1555https://ai.meta.com/blog/meta-llama-3-1/?_fb_noscript=1 (Dubey et al., 2024) even naturally support a context length of 128,000 tokens, which appears promising to process the long patent descriptions. Other LLMs also showed a trend towards long context capabilities, such as Falcon-180B (Almazrouei et al., 2023), Gemini 1.5666https://blog.google/technology/ai/google-gemini-next-generation-model-february-2024/ (Reid et al., 2024) and Claude 3.5777https://www.anthropic.com/news/claude-3-5-sonnet (Anthropic, 2024).
@@ 3.4 LLMs for Patent Processing
Patent texts are distinct from everyday texts with respect to long context length, in-depth technical complexity, and high precision to ensure the patent can be granted, defended, and enforced. This type of language often necessitates specific training and experience in patents for accurate reading and interpretation. However, human readers usually struggle with certain aspects of patent texts, such as ambiguous names, phrases that conflict with their prior understanding, or numerous terms that the description redefines within the context of the specific patent itself. The dense, specialized terminology and unconventional syntactic structures common in patents often pose significant barriers to comprehension, even for those with experience in technical fields. In contrast, LLMs trained on datasets specifically tailored to patent language are theoretically well-equipped to handle these challenges. They are designed to process complex syntactic structures, manage long-range dependencies, and incorporate newly coined terms or domain-specific jargon that diverges from everyday language. By leveraging the specialized vocabulary of experts in the field, these models can navigate the nuanced requirements of patent texts with a level of consistency and scope that may exceed human capabilities. Therefore, the recent advancements in LLMs (Zhao et al., 2023) for generative tasks and language processing appear ideally suited for the unique characteristics of patent literature. They promise large benefits in areas, such as patent drafting, prior art search, and examination.
@@
Despite the apparent fundamental compatibility of LLMs for knowledge extraction and language processing, the application of LLMs in the patent domain remains underdeveloped and not yet highly prominent. Previous studies used word embeddings (e.g., Word2Vec (Mikolov et al., 2013)) and deep learning models (e.g., LSTM (Hochreiter and Schmidhuber, 1997)) for patent analysis tasks. As transformers (Vaswani et al., 2017) showed significant potential in text processing, researchers started to develop transformer-based language models, such as BERT (Devlin et al., 2018) and GPT (Radford et al., 2018). The recent large-sized models with outstanding capabilities have not been extensively investigated in the patent field. Some representative general LLMs that are worth exploring include the Llama-3 family (Dubey et al., 2024), Mistral (Jiang et al., 2023a), Mixtral (Jiang et al., 2024a), GPT-4 (OpenAI, 2023), Claude 3 (Anthropic, 2024), DeepSeek-V3 (Liu et al., 2024), and Gemini 1.5 (Reid et al., 2024). Researchers have also developed patent-specific LLMs, such as PatentGPT-J (Lee, 2023) and PatentGPT (Bai et al., 2024). However, PatentGPT-J has shown limited performance in patent text generation tasks (Jiang et al., 2024c), and PatentGPT, though promising, is a recent development that is not yet publicly available. Significant work related to patent-specific LLMs remains to be done. Moreover, since patents represent a type of legal document, law-specific LLMs are also worth investigating, such as SaulLM (Colombo et al., 2024).
@@
Previous research efforts have included patent analysis, data extraction, and automation of procedures. However, the lack of benchmark tests, such as reference datasets and established metrics, hinders performance evaluation and comparison across different methods. The effectiveness of LLMs depends on the quality of training data. To this end, we have compiled sources and databases of patents with curated datasets tailored for various patent-related tasks in Section 4. Although patent offices have released raw documents for years, publicly available datasets for specific tasks remain scarce. Numerous studies continue to use closed-source data for training and evaluation. Furthermore, patent offices do not provide pre-processed data or broad access to well-structured documents from the process around patents. Although in many countries, patent documents as the disclosure of inventions are considered public domain, they offer only the manual review of individual documents. [^9] This aspect may reflect an important discussion to be led by the field and society. As in other domains, there may be a missing or unclear consensus on which data can be used for machine learning. In stark contrast to other domains, where even books and copyrighted materials may be processed in large numbers, sometimes to create new works that compete with the originals, the patent system is a bit different. The entire modern patent system was developed and is justified as a social contract, trading invention disclosure for temporary exclusion from competition. The descriptions are often public domain and related correspondence with patent offices is made public to ensure transparency. Thus, imposing paywalls would typically violate the mission of patent offices. Admittedly, big-data methods and artificial intelligence were not considerations during the establishment of these legal frameworks. Patent offices may still restrict or complicate larger-scale data access to parts or all contents to avoid commercial exploitation. Some patent offices offer costly paid database access as a stream of revenue, which can be prohibitively expensive for academic research. Important context to know is that in many places patent offices need to fund their own operating expenses (United States Patent and Trademark Office, 2023; President of the European Patent Office and Administrative Council of the European Patent Organisation, 2021; van Pottelsberghe de la Potterie and François, 2009; Frakes and Wasserman, 2014). Therefore, society—which the patent offices are ultimately meant to serve—must reach a consensus on how to manage data access for commercial versus academic purposes, potentially leading to both proprietary and open-source models. The pre-processing steps to formulate structured patent datasets generally involve: segmenting patent documents into clearly defined sections to enhance downstream tasks, such as abstract, claims, description, etc; and rectifying irregularity issues, such as missing fields, erroneous characters, or formatting issues.
@@
3.5 Multimodal Techniques
A patent is not merely a text document but can include drawings, i.e., visual components. Thus, multimodal methods such as CLIP (Radford et al., 2021) and Vision Transformers (Dosovitskiy, 2020) may unlock this potential. Multimodal methods in patent processing integrate diverse data types, such as text, images, and quantitative information, to enhance tasks such as classification and retrieval. The combination of their complementary strengths of different modalities may lead to more comprehensive and accurate results. Lee et al. (2022) for instance introduced a multimodal deep-learning model that combined textual content with quantitative patent information, which resulted in improved performance for patent classification tasks. Additionally, multimodal approaches that integrate textual descriptions with visual content have shown promise in enhancing patent retrieval (Lo et al., 2024). Furthermore, Lin et al. (2023) proposed multimodal methods to extract structural and visual features to effectively measure patent similarity. Such multimodal similarity detection may improve the efficiency of patent examination. However, as illustrated in Fig. 3, many drawings in patents are generic without the corresponding description, and some drawings suffer from poor resolution, pixelation, or low quality. This may be one of the reasons why current multimodal methods in patent processing are scarce. Moreover, textual elements, especially the claims and descriptions, are the primary carriers of legally binding information in patents, particularly features. While drawings provide valuable support and document embodiments, the text is essential for defining the scope, novelty, and context of each invention. Consequently, the survey particularly examines NLP approaches and includes representative multimodal methods where available.